OpenStudy (chairbear101k12):

Evaluate. 2^5/2^3 A) 1 B) 2/8 C) 4 D) 8

OpenStudy (vijeya3):

Are ya aware with the laws of exponentiation?

OpenStudy (welshfella):

when you divide these terms you subtract the exponents

rishavraj (rishavraj):

\[\frac{ a^x }{ a^y } = a^{x-y}\]

OpenStudy (chairbear101k12):

i just need the answer really bad if yall could just tell me that would be amazing

OpenStudy (vijeya3):

LOL that would be 2^2

OpenStudy (chairbear101k12):

@Vijeya3 that is not a amswer choice

rishavraj (rishavraj):

@Vijeya3 u aint supposed to give answers right away :P

OpenStudy (mathmale):

Unfortunately, OpenStudy was not set up to provide answers with no effort on your part. There are definite and clear rules for dealing with exponents, and you need to know and be able to apply them.

OpenStudy (vijeya3):

Oh yeah sorry That would evaluate to 4 so the right answer is C :)

OpenStudy (mathmale):

rishavraj has already given you the formula you need. Apply it here, please.

OpenStudy (vijeya3):

@rishavraj I was gonna go the right way....the person wants just direct answers...so i helped him the way he desired

OpenStudy (mathmale):

You are evaluating\[\frac{ 2^5 }{ 2^3 }\]

OpenStudy (mathmale):

Apply rishavraj's formula.

OpenStudy (mathmale):

Bear: Glad to help, IF and ONLY IF you first show some effort of your own to solve this problem.